# Eigenvalue analysis of three-state quantum walks with general coin matrices

### v4

@@ IV.1 One-defect model
$C_{x}=\begin{cases}e^{i\theta}F&x=0,\\ F&x\neq 0.\end{cases}$
@@ IV.2 Two-phase model
$C_{x}=\begin{cases}F&x<0,\\ e^{i\theta}F&x\geq 0.\end{cases}$
@@ V Conclusion and Discussion
In this paper, we developed a general transfer matrix method for analyzing the spectral properties of space-inhomogeneous three-state quantum walks (QWs) on the integer lattice with general coin matrices. This extends previous results [26], which were limited to specific classes such as Grover-type coins, to a broader class of models including non-Grover-type coins like the Fourier matrix. Our approach reduces the eigenvalue problem of the original three-state QW to a tractable two-component recurrence relation, enabling an explicit characterization of eigenvalues and eigenstates through a $2\times 2$ transfer matrix formalism. We derived a necessary and sufficient condition for the eigenvalue problem in the case of two-phase QWs with a finite number of defects, providing a mathematical characterization and enabling a precise analysis of localization phenomena in these systems.
@@
These results not only provide new insights into the spectral theory of three-state QWs but also highlight the effectiveness of our method in studying localization beyond symmetric models such as the Grover walk. Our framework is particularly relevant for exploring quantum transport phenomena, topological phases, and the stability of localized states under perturbations. Moreover, it opens up possibilities for extending similar spectral analyses to multi-state QWs involving many particles [49, 50, 51], as well as to higher-dimensional QWs [52, 53].
@@
Therefore, a future work could be to extend our framework to multi-particle, multi-state QWs and to higher-dimensional lattices. For a two-dimensional grid [52] that remains translationally invariant along one axis but not the other, we can apply a Fourier transform in the invariant direction, reducing the problem to an effective four-state walk on a one-dimensional lattice with two self-loops. This reduction lets us deploy our transfer-matrix formalism unchanged, opening the door to rigorous, analytic treatment of two-dimensional QWs, precisely the regime where rich localization phenomena, non-trivial topological phases, and scalable quantum-algorithmic applications are expected to emerge.
@@
Data availability: The simulation code used in this work is available online at https://github.com/CKiumi/quantum_walk.git.
@@ Acknowledgment
The author C. Kiumi was supported by JSPS KAKENHI Grant Number JP22KJ1408 and JST ASPIRE Japan Grant Number JPMJAP2319.
@@ Appendix A Proof of Proposition 1
so that $\iota^{-1}\tilde{\Psi}\in\mathcal{H}$ satisfies the eigenvalue equation $U(\iota^{-1}\tilde{\Psi})=e^{i\lambda}(\iota^{-1}\tilde{\Psi})$. Then, the transfer matrix is derived by the reformulation of Eqs. 8 and 9 as
@@
From this expression, the transfer matrix can be constructed only if $A_{x}(\lambda)\neq 0$, and we will discuss this case later in this section. Recall that the coin matrix $C_{x}$ is a $3\times 3$ unitary matrices, which is written as below:
$C_{x}=\left[\begin{array}[]{ c c c }a_{x}^{(1,1)}&a_{x}^{(1,2)}&a_{x}^{(1,3)}\\ a_{x}^{(2,1)}&a_{x}^{(2,2)}&a_{x}^{(2,3)}\\ a_{x}^{(3,1)}&a_{x}^{(3,2)}&a_{x}^{(3,3)}\end{array}\right].$
@@
$C_{x}=e^{i\Delta_{x}}\left[\begin{array}[]{ c c c }\overline{\det C_{x}^{(1,1)}}&-\overline{\det C_{x}^{(1,2)}}&\overline{\det C_{x}^{(1,3)}}\\ -\overline{\det C_{x}^{(2,1)}}&\overline{\det C_{x}^{(2,2)}}&-\overline{\det C_{x}^{(2,3)}}\\ \overline{\det C_{x}^{(3,1)}}&-\overline{\det C_{x}^{(3,2)}}&\overline{\det C_{x}^{(3,3)}}\end{array}\right],$
where $e^{i\Delta_{x}}=\det C_{x}$ and $C_{x}^{(i,j)}$ denote the $2\times 2$ submatrix obtained by excluding the $i$-th row and $j$-th column of $C_{x}$. Using the above expression, we can simplify each component of the transfer matrix as:
@@
$=\frac{\det C_{x}^{(2,2)}e^{i\lambda}-e^{i\Delta_{x}}}{e^{i\lambda}-a_{x}^{(2,2)}}=-e^{i(\lambda+\Delta_{x})}\frac{e^{-i\lambda}-\overline{a_{x}^{(2,2)}}}{e^{i\lambda}-a_{x}^{(2,2)}}.$
@@
Nextly, we assume $A_{x}(\lambda)\neq 0$ where we cannot construct the transfer matrix. First we state the following lemma.
@@
$B_{x}(\lambda)=\frac{a_{x}^{(1,3)}\overline{a_{x}^{(3,3)}}+a_{x}^{(1,1)}\overline{a_{x}^{(3,1)}}}{\overline{a_{x}^{(3,3)}}-e^{-i\Delta_{x}}a_{x}^{(1,1)}a_{x}^{(2,2)}}=e^{i\Delta_{x}}\frac{\overline{a_{x}^{(3,2)}}}{a_{x}^{(2,1)}},$
@@ Appendix B Proof of Main results
Theorem 1 and its corollaries, which establish the necessary and sufficient conditions for the eigenvalue problem, and they can be directly proved by following the discussion in Section III.3 to consider the square summability of $\tilde{\Psi}$ in Proposition 1. Lemma 2 and Lemma 3 are the core properties of the transfer matrix, which are essential for the proof of the main theorem. The proof of these propositions is provided in Section B.1 and Section B.2, respectively.
@@ B.2 Proof of Lemma 3
Assuming $\mathrm{tr}(T_{\pm\infty})\neq 0$, by (1), we have
@@
Substituting this into (2) gives
@@
Thus, the condition $|\mathrm{tr}(T_{\pm\infty})|^{2}-4\leq 0$ is necessary and sufficient for (2) to hold, meaning $\left|\zeta_{\pm\infty}^{+}\right|=\left|\zeta_{\pm\infty}^{-}\right|$. If $\mathrm{tr}(T_{\pm\infty})=0$, it is clear that $\left|\zeta_{\pm\infty}^{+}\right|=\left|\zeta_{\pm\infty}^{-}\right|$ holds as well, with $|\mathrm{tr}(T_{\pm\infty})|^{2}-4=-4\leq 0$.

### v5

@@ IV.1 One-defect model
$\mathsf{C}_{x}=\begin{cases}e^{i\theta}F&x=0,\\ F&x\neq 0.\end{cases}$
@@ IV.2 Two-phase model
$\mathsf{C}_{x}=\begin{cases}F&x<0,\\ e^{i\theta}F&x\geq 0.\end{cases}$
@@ V Conclusion and Discussion
In this paper, we developed a general transfer matrix method for analyzing the spectral properties of space-inhomogeneous three-state quantum walks (QWs) on the integer lattice with general coin matrices. This extends previous results [26], which were limited to specific classes such as Grover-type coins, to a broader class of models including non-Grover-type coins like the Fourier matrix. Furthermore, we explicitly characterized the essential spectrum, identifying both the flat-band eigenvalues with infinite multiplicity and the absolutely continuous spectrum. This completes the full spectral classification of the model. Our approach reduces the eigenvalue problem of the original three-state QW to a tractable two-component recurrence relation, enabling an explicit characterization of eigenvalues and eigenstates through a $2\times 2$ transfer matrix formalism. We derived a necessary and sufficient condition for the eigenvalue problem in the case of two-phase QWs with a finite number of defects, providing a mathematical characterization and enabling a precise analysis of localization phenomena in these systems.
@@
These results not only provide new insights into the spectral theory of three-state QWs but also highlight the effectiveness of our method in studying localization beyond symmetric models such as the Grover walk. Our framework is particularly relevant for exploring quantum transport phenomena, topological phases, and the stability of localized states under perturbations.
V.1 Physical implications
Our transfer-matrix framework provides a complementary and rigorous method for analyzing the spectrum of three-state quantum walks. In particular, the eigenvalue conditions established in Theorem 1 allow for a precise characterization of localized states induced by defects or phase boundaries, while the resulting three-way spectral decomposition offers a clear physical interpretation of the dynamics generated by general three-state coins:
  • Discrete eigenvalues in $\sigma_{\mathrm{disc}}(U)$, correspond to genuine defect or edge modes that decay exponentially away from the inhomogeneous region. Any initial state with nonzero overlap on such eigenvectors exhibits persistent localization near the defect, as confirmed numerically in Section IV.
  • Flat-band points in $\sigma^{\infty}_{\mathrm{ess}}(U)$ arise from special symmetries of the tail coins and support massively degenerate, compactly supported states. These modes are spatially localized but form part of the essential spectrum due to their infinite multiplicity.
  • The absolutely continuous part $\sigma_{\mathrm{ac}}(U)$ corresponds to extended states. For initial conditions overlapping only with this part, the walk exhibits ballistic spreading and delocalization.
Altogether, the transfer-matrix approach not only pinpoints the eigenvalues and their associated eigenstates, but also integrates them into a unified spectral picture in which extended bands, possible flat bands, and defect-induced localized modes coexist. This structure explains the diverse localization and spreading phenomena observed in space-inhomogeneous three-state quantum walks and highlights the utility of spectral methods for their analysis.
V.2 Connections to topological phases and quantum algorithms
From a physical perspective, the exponentially localized eigenstates that appear at boundaries between different asymptotic coins can be naturally interpreted as edge states in the language of topological quantum walks [1, 2]. In particular, in models where the bulk time-evolution operators admit a topological classification, the transfer-matrix method detects discrete eigenvalues in quasi-energy gaps as elements of $\sigma_{\mathrm{disc}}(U)$ in Theorem 2; these eigenvalues correspond to the gap states associated with bulk–edge correspondence. On a more fundamental level, it would be an interesting direction to explore the relationship between the spectral structure analyzed in this study and the index theorems [54]. Specifically, clarifying whether the number of solutions to $\chi(\lambda)=0$ corresponds to a certain Fredholm index would provide a more rigorous mathematical foundation for the bulk–edge correspondence in three-state models. Moreover, the structure of the flat-band sector carries additional physical significance: gapped flat bands are relatively easy to isolate and can support strongly correlated or Mott phase, whereas gapless flat bands may give rise to critical or semimetal-like behavior and are associated with unconventional transport phenomena [55]. A systematic derivation of the relevant bulk invariants for general three-state walks is left for future study.
Quantum walks also provide a powerful framework for designing search algorithms on graphs and lattices, and more generally, preparing target states with high probability. A key insight is that the success probability and time-to-solution of quantum walk search depend critically on the spectral properties of the underlying walk operator. For instance, lackadaisical quantum walks [56, 57, 58, 59, 60] where additional self-loops are introduced at each vertex—have been widely studied as an effective mechanism to improve search efficiency on both regular lattices and more general graphs. [9] also showed that by introducing an interpolated walk—a modification that effectively adds a self-loop like stay transition to the underlying chain—one can construct a general framework for quantum-walk search algorithms on arbitrary graphs. In these models, the presence of self-loops modifies the coin operator and hence the spectral gap, directly influencing the algorithmic speedup.
The connection between quantum singular value transformation (QSVT)—a general framework for quantum algorithms and split-step quantum walks used to probe topological phases has recently been explored in [21]. Because QSVT is fundamentally about shaping and controlling the spectrum of embedded operators within a unitary evolution, it is natural to speculate that similar spectral principles may underlie both algorithmic speedups and topological phenomena. In particular, introducing suitable symmetries or structured coins in our three-state model may enrich its spectral landscape in ways reminiscent of topological quantum walks. Establishing a unified viewpoint that bridges topological phases and quantum speed-up via the spectral analysis of quantum walks therefore presents an intriguing direction for future research.
V.3 Limitation and future works
While our analysis covers a broad class of space-inhomogeneous three-state quantum walks, it still has important limitations. In this work we restrict attention to walks that are asymptotically homogeneous on both sides (two-phase tails) with only finitely many defects in between. This setting already includes many standard models such as finite-defect and two-phase interfaces, but it does not encompass completely general inhomogeneous coins with arbitrarily long-range or disordered variations.
A natural direction for future work is to extend the transfer-matrix formalism beyond the two-phase setting. One promising avenue is to treat walks periodic tails by directly applying techiniques in [48]. An even richer generalization is to incorporate random or quasi-periodic inhomogeneities, where the spectral properties can be governed by Lyapunov exponents and ergodic cocycle theory, following approaches used in studies of localization and fractality in inhomogeneous quantum walks [61], and in recent works on Anderson localization for quasi-periodic CMV matrices and quantum walks [62, 63].
@@ V.3 Limitation and future works
Moreover, it opens up possibilities for extending similar spectral analyses to multi-state QWs involving many particles [64, 65, 66], as well as to higher-dimensional QWs [67, 68]. A future work could be to extend our framework to multi-particle, multi-state QWs and to higher-dimensional lattices. For a two-dimensional grid [67] that remains translationally invariant along one axis but not the other, we can apply a Fourier transform in the invariant direction, reducing the problem to an effective four-state walk on a one-dimensional lattice with two self-loops. This reduction lets us deploy our transfer-matrix formalism unchanged, opening the door to rigorous, analytic treatment of two-dimensional QWs, precisely the regime where rich localization phenomena, non-trivial topological phases, and scalable quantum-algorithmic applications are expected to emerge.
We expect that this research will provide a foundational basis for the rigorous spectral analysis of the broader and more complex quantum-walk models outlined above, ultimately contributing to a comprehensive mathematical theory of inhomogeneous quantum walks.
@@
Data availability: The simulation code used in this work is available online at [69].
@@ Acknowledgment
This work is supported by JSPS KAKENHI, Grant Number JP22KJ1408, JST ASPIRE Japan, Grant Number JPMJAP2319 and JST PRESTO Japan, Grant Number JPMJPR25F1.
@@ Appendix A Proof of Proposition 1
[…]
@@
From this expression, the transfer matrix can be constructed only if $A_{x}(\lambda)\neq 0$, and we will discuss this case later in this section. Recall that the coin matrix $\mathsf{C}_{x}$ is a $3\times 3$ unitary matrices, which is written as below:
$\mathsf{C}_{x}=\left[\begin{array}[]{ c c c }a_{x}^{(1,1)}&a_{x}^{(1,2)}&a_{x}^{(1,3)}\\ a_{x}^{(2,1)}&a_{x}^{(2,2)}&a_{x}^{(2,3)}\\ a_{x}^{(3,1)}&a_{x}^{(3,2)}&a_{x}^{(3,3)}\end{array}\right].$
@@
$\mathsf{C}_{x}=e^{i\Delta_{x}}\left[\begin{array}[]{ c c c }\overline{\det\mathsf{C}_{x}^{(1,1)}}&-\overline{\det\mathsf{C}_{x}^{(1,2)}}&\overline{\det\mathsf{C}_{x}^{(1,3)}}\\ -\overline{\det\mathsf{C}_{x}^{(2,1)}}&\overline{\det\mathsf{C}_{x}^{(2,2)}}&-\overline{\det\mathsf{C}_{x}^{(2,3)}}\\ \overline{\det\mathsf{C}_{x}^{(3,1)}}&-\overline{\det\mathsf{C}_{x}^{(3,2)}}&\overline{\det\mathsf{C}_{x}^{(3,3)}}\end{array}\right],$
where $e^{i\Delta_{x}}=\det\mathsf{C}_{x}$ and $\mathsf{C}_{x}^{(i,j)}$ denote the $2\times 2$ submatrix obtained by excluding the $i$-th row and $j$-th column of $\mathsf{C}_{x}$. Using the above expression, we can simplify each component of the transfer matrix as:
@@
$=\frac{\det\mathsf{C}_{x}^{(2,2)}e^{i\lambda}-e^{i\Delta_{x}}}{e^{i\lambda}-a_{x}^{(2,2)}}=-e^{i(\lambda+\Delta_{x})}\frac{e^{-i\lambda}-\overline{a_{x}^{(2,2)}}}{e^{i\lambda}-a_{x}^{(2,2)}}.$
@@
Nextly, we assume $A_{x}(\lambda)=0$ where we cannot construct the transfer matrix. First we state the following lemma.
@@
$B_{x}(\lambda)=\frac{a_{x}^{(1,3)}\overline{a_{x}^{(3,3)}}+a_{x}^{(1,1)}\overline{a_{x}^{(3,1)}}}{\overline{a_{x}^{(3,3)}}-e^{-i\Delta_{x}}a_{x}^{(1,1)}a_{x}^{(2,2)}}=e^{i\Delta_{x}}\frac{\overline{a_{x}^{(3,2)}}}{a_{x}^{(2,1)}},\$
@@ Appendix B Proof of Main results
Theorem 1 and its corollaries, which establish the necessary and sufficient conditions for the eigenvalue problem, and they can be directly proved by following the discussion in Section III.1 to consider the square summability of $\tilde{\Psi}$ in Proposition 1. Lemma 2 and Lemma 3 are the core properties of the transfer matrix, which are essential for the proof of the main theorem. The proof of these propositions is provided in Section B.1 and Section B.2, respectively.
@@ B.2 Proof of Lemma 3
Assuming $\mathrm{tr}(T_{\pm\infty})\neq 0$, by Eq. 14, we have
@@
Substituting this into Eq. 15 gives
@@
Thus, the condition $|\mathrm{tr}(T_{\pm\infty})|^{2}-4\leq 0$ is necessary and sufficient for Eq. 15 to hold, meaning $\left|\zeta_{\pm\infty}^{+}\right|=\left|\zeta_{\pm\infty}^{-}\right|$. If $\mathrm{tr}(T_{\pm\infty})=0$, it is clear that $\left|\zeta_{\pm\infty}^{+}\right|=\left|\zeta_{\pm\infty}^{-}\right|$ holds as well, with $|\mathrm{tr}(T_{\pm\infty})|^{2}-4=-4\leq 0$.